OpenStudy (anonymous):

what is the energy required or released to form a dipole from a neutral polarizable molecule.

OpenStudy (turingtest):

The potential energy of a magnetic dipole in a magnetic field is\[U=-\vec\mu\cdot\vec B\]an electric dipole has the same formula for potential energy that any two charges have\[U=\frac1{4\pi\epsilon_0}\frac{q_1q_2}r\]

OpenStudy (vincent-lyon.fr):

Energy of the two opposite charges will be: \(U=-\Large \frac1{4\pi\epsilon_0}\frac{q^2}{a}=-\frac1{4\pi\epsilon_0}\frac{P^2}{a^3}\) if distance between the charges is a and P=qa

OpenStudy (anonymous):

Doesn't that consider potential to be zero at infinity? What we are doing is separating charge from zero to distance r right? We are not bringing charge from infinity to r distance. \[ \int_r^0 {1 \over 4 \pi \epsilon _ 0}{ q^2 \over r^2} = \left [ - kq^2 \over r\right ]_0^r = kq^2 /0 - kq^2/r\]

OpenStudy (vincent-lyon.fr):

Well, the question was : "What is the energy stored in the dipole" which means exactly, "what work can you obtain when the charges are taken to infinity with zero-velocity".

OpenStudy (anonymous):

sorry ... my mistake!! any idea on what i mean??

OpenStudy (anonymous):

please recheck edited question